Q1) Suppose a firm has 33.30 million shares of common stock outstanding at a price of $26.25 per share. The firm also has 361000.00 bonds outstanding with a current price of $1,108.00. The outstanding bonds have yield to maturity 6.63%. The firm's common stock beta is 1.811 and the corporate tax rate is 40.00%. The expected market return is 11.04% and the T-bill rate is 3.42%. Compute the following: a) Weight of Equity of the firm (2 points) b) Weight of Debt of the firm (2 points) c) Cost of Equity of the firm (2 points) d) After Tax Cost of Debt of the firm (2 points) e) WACC for the Firm (2 points)
Answer #1

Value of Equity = 33,300,000(26.25)

Value of Equity = $874,125,000

Value of Debt = 361,000(1,108)

Value of Debt = $399,988,000

Weight of Equity = 874,125,000/(874,125,000 + 399,988,000)

Weight of Equity = 0.6861

Weight of Debt = 399,988,000/(874,125,000 + 399,988,000)

Weight of Debt = 0.3139

Calculating After tax Cost of Debt,

After-tax cost of Debt = (1 - 0.40)(0.0663)

After-tax cost of Debt = 3.98%

Calculating Cost of Equity,

Cost of Equity = 0.0342 + 1.811(0.1104 - 0.0342)

Cost of Equity = 17.22%

WACC = 0.3139(0.0398) + 0.6861(0.1722)

WACC = 13.06%